# Uncovering LLM-Generated Code: A Zero-Shot Synthetic Code Detector via Code Rewriting

### v1

@@ 3.2. Overview
Our method is based on the intuition that every programmer has a unique coding style and tends to solve coding problems following their own routines and habits. Similarly, generative models can be likened to senior programmers, each with their distinctive coding style. This method is inspired by the code reviewing and refactoring process. In these stages, the original author of the code usually does not participate. Managers assign these tasks to other programmers to prevent the original authors from being stuck in their habitual routines, which might hinder them from making significant improvements. In a similar vein, Large Language Models tend to exhibit consistent code writing patterns due to inherent biases in their training data. Therefore, LLMs can be viewed as programmers with unique coding styles and habitual routines shaped by the common biases present in their training data.
Therefore, we hypothesize that when asking generative models to rewrite synthetic code, the differences between the rewritten and the original code tend to be smaller than when the original code is human written. Refer to Figure 2 for an intuitive understanding. Based on this hypothesis, we propose a zero-shot synthetic code detection method using Code Rewriting and Similarity Measurement. The overall design principle is: for a given code snippet $x$, we generate code rewriting multiple times using a generative LLM $\mathcal{G}$. Then we use a code similarity model $\mathcal{M}$ to measure the average similarity between rewritten code and original code $x$. We train a code similarity model $\mathcal{M}$ using self-supervised contrastive learning (Gao et al., 2021) on unlabeled code to measure the code similarity effectively. Consequently, we use the average similarity score as the detection metric. We summarize the proposed approach in Algorithm 1. Next, the two parts of Code Rewriting and Similarity Measurement will be explained in detail.
@@ 3.3. Code Rewriting
Given a code snippet $x$, we generate a rewriting of $x$ utilizing the chain of thought prompting method (Wei et al., 2022b). We prompt the generative model $\mathcal{G}$ with the original code snippet $x$ and instruct the model to initially generate an analysis and explanation of $x$, followed by a potential rewrite $x^{\prime}$. The intermediate code analysis can help the model to understand the original code and generate a valid rewrite according to the analysis. It should be noted that $x$ is normalized by omitting all comments and empty lines. Because comments are very easy to manipulate while maintaining their functionality. They can either be a source of noise or be utilized to evade detection.
@@ 3.4. Similarity Measurement
To effectively measure the similarity between the rewritten code and the original code, we require a code similarity model $\mathcal{M}$, which can predict a similarity score $S$ for the input code pair. In the domain of AI for software engineering, code similarity learning is an activate area of research. Many studies address this problem by focusing on code representation learning. This involves generating dense semantic representations by leveraging various code structures (Feng et al., 2020; Ye et al., 2021; Wang et al., 2021b; Guo et al., 2021; Wang et al., 2022; Li et al., 2022c) and using vector distances to quantify similarity. We use GraphCodeBERT (Guo et al., 2021) as the similarity model since its model checkpoint is off-the-shelf, widely used and has been pre-trained on large-scale code.
To obtain better function-level representation from a pre-trained transformer, we employ self-supervised contrastive learning to enhance the sentence-level representation of GraphCodeBERT further. Following (Gao et al., 2021), we adapt the unsupervised SimCSE method to the code domain, using standard dropout as the data augmentation method for contrastive learning. For an input code snippet $x$, we use the last-layer hidden states of [CLS] token as code representation. During the unsupervised training stage, we add an MLP layer to obtain the final representation $\mathbf{h}$ of $x$.
@@
Where $\tau$ is a temperature hyperparameter and is set to $0.1$. The $sim(\mathbf{h}_{i},\mathbf{h}^{\prime}_{i})$ is the cosine similarity $\frac{\mathbf{h}_{i}\mathbf{h}^{\prime}_{i}}{|\mathbf{h}_{i}||\mathbf{h}^{\prime}_{i}|}$.
The MLP layer is dropped after the unsupervised contrastive learning stage, and we only use the last-layer hidden states of the [CLS] token as snippet representation following (Gao et al., 2021). For a given code snippet $x$ and its rewriting $x^{\prime}$ sampled from $\mathcal{G}$ as introduced in Section 3.3, we obtain the final embedding of $x$ and $x^{\prime}$ by feeding them to the similarity model $\mathcal{M}$. It is worth noting that our framework is not dependent on a specific similarity model. $\mathcal{M}$ can be any other implementations that can accurately model the code similarity, such as OpenAI’s text embedding services (Greene et al., 2022).
@@
A larger $m$ can result in a more accurate expectation estimation but at the cost of generating more rewritten code. Our experiments found that using only 4 rewrites can achieve outstanding detection performance.
@@ 4.1. Benchmarks
Considering there is no existing benchmark for evaluating synthetic code detectors. We build two synthetic code detection benchmarks in Python using APPS (Hendrycks et al., 2021) and MBPP (Austin et al., 2021). The choice of Python as the programming language for constructing our benchmark stems from several key considerations. Firstly, a significant portion of training data for prevalent code generation models like CodeGen (Nijkamp et al., 2022), Incoder (Fried et al., 2022), and StarCoder (Li et al., 2023) is constituted by Python code. These models exhibit notably superior code generation capabilities in Python than other languages. Consequently, the challenge and significance of detecting code generated by LLMs are most pronounced within the context of Python. Secondly, popular datasets used for code generation evaluation, such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and APPS (Hendrycks et al., 2021), primarily provide Python test cases. We adopted Python as our chosen language to align with this prevailing practice. Furthermore, Python ranks among the most widely used programming languages, adding practical significance to validating our detection performance. Specifically, for the generation of synthetic code, we utilize the natural language description of each sample as a prompt and employ four widely-used code generation tools (CodeLlama (Rozière et al., 2023), StarChat (Tunstall et al., 2023), GPT-3.5 and GPT-4 (OpenAI, 2023)).
@@
We selected two open-sourced Code Large Language Models, CodeLlama-13B-Instruct (Rozière et al., 2023) and StarChat-Alpha (Tunstall et al., 2023), along with two proprietary generation APIs, GPT-3.5-turbo and GPT-4 (OpenAI, 2023). According to the survey (StackOverflow, 2023) conducted by StackOverflow, GPT-3.5-turbo is the most extensively used tool in daily programming. GPT-4 represents the state-of-the-art in code generation tools to date. Consequently, we have chosen these two models to typify proprietary generation APIs. CodeLlama (Rozière et al., 2023) and StarChat (Tunstall et al., 2023) are two potent open-sourced code generation tools, with their 34B versions exhibiting competitive performance when evaluated against GPT-3.5-turbo. We opted for the 13B and 15.5B versions due to resource constraints. Nonetheless, these versions also demonstrate superior generation performance.
@@
APPS is a benchmark for code generation. Each sample in APPS is a coding problem taken from coding contest websites, e.g., CodeForces, LeetCode, and HackerRank, and contains problem descriptions and submitted solutions for the problem. The original APPS Benchmark comprises 5,000 test data and 5,000 training data instances. In our construction, we exclusively utilized the test data. The rationale for excluding the training data stems from concerns that the training data might have been incorporated into the training data of code LLMs. If the training data are using as prompts and those prompts are present in the LLM’s training data, it could potentially result in code generation that perfectly matches human-written code, leading to inaccuracies in evaluating the detection outcomes. Within the pool of 5,000 test data instances, we identified and removed prompts with web hyperlinks from the markdown format of coding challenges, as such links could affect code quality and are rarely used by developers in practice. This reduced the dataset to 3,846 instances. We then randomly selected 1,540 instances (40%) from this refined set and used 770 of them for synthetic code generation.
For those 770 problems, we generate synthetic code using the chain of thought prompting (Wei et al., 2022b) to improve the generation quality using four generation tools. Figure 4 shows the exact prompt used for code generation. The generation temperature is set to $0.7$, and $top\_p$ is set to 0.95. We removed all the comments and empty lines for the generated code. For the remaining 770 problems (1540 minus 770), we randomly sample a solution from all valid solutions as human-written code. The detailed benchmark statistics are listed in Table 1.
@@
Table 1. APPS Benchmark Statistics. We count the averated number of chars and lines for both human-written/synthetic code for all versions of benchmark generated by four different code generation tools.
@@
MBPP is another Benchmark for evaluating AI code generation. It contains 1,000 crowd-sourced Python programming problems designed to be solvable by entry-level programmers, covering programming fundamentals, standard library functionality, etc. Compared to APPS, MBPP’s problems are more straightforward and frequently encountered in daily programming practice. We follow the same construction process as APPS. The detailed benchmark statistics are listed in Table 2.
@@ 4.2. Similarity Model Training Details
We continue to train a code similarity model by exploiting unsupervised SimCSE from initial GraphCodeBERT as introduced in Section 3.4. In the SimCSE training stage, we collect 160k code snippets from CodeSearchNet (Husain et al., 2020) Python subset as training data. The GraphCodeBERT model trained by SimCSE is fixed for all subsequent experiments once training is complete. We use the Adam optimizer and set the maximum learning rate to $41e-$ 4 with linear decay. We use 4 NVIDIA 3090s with batch size 16 on each GPU and train the model for 5 epochs. In addition to GraphCodeBERT model, we also experimented with the UniXcoder (Guo et al., 2022) as another similarity model by following the same SimCSE training process.
@@ 4.3. Detector LLMs
Our approach necessitates a rewriting LLM, $\mathcal{G}$, to function as the detector LLM. Additionally, other zero-shot detectors for text, such as GLTR and DetectGPT, require a scoring LLM for likelihood estimation, as introduced in Section 2.2. We examine CodeLlama, StarChat-Alpha, and GPT-3.5 as detector LLMs for our method, while using CodeLlama and StarChat-Alpha for baseline comparison. We employ nucleus sampling for code rewriting, setting the top-$p$ to $0.95$ and the temperature to $0.8$.
@@ 4.5. Baselines
We consider the following zero-shot detection methods using a surrogate model to approximate the true distribution of the generative model: $\mathbf{\log p(x)}$ (Gehrmann et al., 2019), LogRank, Rank, Entropy (Mitchell et al., 2023), LRR, NPR (Su et al., 2023) and DetectGPT (Mitchell et al., 2023). To adapt these detection metrics to code content and effectively approximate the true distribution of the generative code LLM, the surrogate score model should be replaced with LLMs trained on large-scale code content. In this study, we examine two open-sourced code LLMs: StarChat-Alpha and CodeLlama. We also select two supervised detectors and compare them to the zero-shot detectors: GPTZero (GPTZero, 2023) is a leading AI content detection servicen and is trained on millions of synthetic texts sampled from various generative models, including ChatGPT, GPT-4 and Bard. OpenAI-Detector (OpenAI, 2019), an open-sourced detector, is trained on texts sampled from GPT-2 and is based on Roberta-large (Liu et al., 2019).
@@
DetectGPT (Mitchell et al., 2023) use the average token probability disparity between the given text and perturbed texts to detect whether the given text is located at the probability curvature of the generative model. To better adapt DetectGPT for the code domain, we made minor modification. Specifically, we replace the original perturbation model, T5-large (Raffel et al., 2020) with CodeT5-large (Wang et al., 2021a). All other settings are the same as in the original implementation.
@@ 5. Experiment Results
Table 3. Main Results. The first two rows list AUROC scores of two detectors from GPTZero (2023) and OpenAI (2019) across two benchmarks and four Generation Tools. The following three sections detail the AUROC scores of our methods, with variations in $m$, along with seven other zero-shot detectors, using three Detector LLMs. CodeLlama is an abbreviation for CodeLlama-13B-Instruct, StarChat denotes StarChat-Alpha-15B, and GPT-3.5 refers to GPT-3.5-turbo. Red indicates the best performance, while blue indicates the best performance among all seven baselines.
@@ 5.1. Main Results
We list the AUROC score of all zero-shot detectors on the APPS and MBPP benchmarks in Table 3. Among the seven baselines, LogRank, LRR and NPR are the most compelling methods. However, the detection performance of baseline zero-shot detectors, including $\log p(x)$, DetecGPT, LRR and NPR drop significantly on the code benchmarks compared to their reported performance on detecting general texts (Mitchell et al., 2023; Su et al., 2023). This observation holds true regardless of the generation tools and detector LLMs used. The main reason lies in the fact that these SOTA zero-shot text detector are based on the log probability of tokens statistically. In the code domain, code exhibits a more uniform grammatical structure, often with a relatively fixed space for the next code tokens. As a result, there are many ”low-entropy” tokens in code, which significantly reduces the number of effective tokens that SOTA text detection methods can use to differentiate between human-written code and model-generated code, as illustrated in Figure 1. Furthermore, We observe that the previous baselines demonstrate improved performance when the Detector LLM and the Generation LLM are identical, and exhibit a decline in performance when they do not match.
As shown in Table 3, our proposed methods significantly outperform previous methods by all four prevalent code generation tools, with a 20.5% improvement on the APPS benchmark and 29.1% improvement on the MBPP benchmark. This improvement is primarily due to our proposed methods detecting synthetic code from a more holistic perspective instead of relying on token-wise scores. The holistic code rewriting and similarity measurement approach effectively avoids the inherent shortcomings of previous methods when applied to the code domain, while also leveraging the tendency of LLMs to select the most confident and likelihood code during generation.

### v2

@@ 3.2. Overview
Our method is grounded in the intuition that every programmer has a unique coding style and tends to solve coding problems following their own routines and habits. Similarly, generative models can be likened to senior programmers with their distinctive coding style. This method is inspired by the code reviewing and refactoring process. During these stages, the original author of the code typically does not participate. Managers assign these tasks to other programmers to prevent the original authors from being stuck in their habitual routines, which might hinder them from making significant improvements. In a similar vein, LLMs tend to exhibit consistent code writing patterns due to inherent biases in their training data. Therefore, LLMs can be viewed as programmers with unique coding styles and habitual routines shaped by the common biases present in their training data.
Therefore, based on consistent code writing patterns, we hypothesize that when generative LLMs are tasked to rewrite synthetic code, the differences between the rewritten and the original code tend to be smaller compared to when the original code is human-written. Refer to Figure 2 for an intuitive understanding. Building on this hypothesis, we propose a zero-shot synthetic code detection method using Code Rewriting and Similarity Measurement. The overall design principle is as follows: for a given code snippet $x$, we process code rewriting multiple times using a generative LLM $\mathcal{G}$. Then, we use a code similarity model $\mathcal{M}$ to measure the average similarity between rewritten code and original code $x$. We train a code similarity model $\mathcal{M}$ using self-supervised contrastive learning (Gao et al., 2021) on unlabeled code to measure the code similarity effectively. Consequently, we use the average similarity score as the detection metric. We summarize the proposed approach in Algorithm 1. Next, the two parts of Code Rewriting and Similarity Measurement will be explained in detail.
@@ 3.3. Code Rewriting
Given a code snippet $x$, we generate a rewriting of $x$ utilizing the chain of thought prompting method (Wei et al., 2022b). We prompt the generative model $\mathcal{G}$ with the original code snippet $x$ and instruct the model to initially generate an analysis and explanation of $x$, followed by a potential rewrite $x^{\prime}$. The intermediate code analysis can help the model to understand the original code and generate a valid rewrite according to the analysis. It should be noted that $x$ is normalized by omitting all comments and empty lines because comments are very easy to manipulate while maintaining their functionality. They can either be a source of noise or be utilized to evade detection.
@@ 3.4. Similarity Measurement
To effectively measure the similarity between the rewritten code and the original code, we require a code similarity model $\mathcal{M}$, which can predict a similarity score $S$ for the input code pair. In the domain of AI for software engineering, code similarity learning is an active area of research. Many studies address this problem by focusing on code representation learning. This involves generating dense semantic representations by leveraging various code structures (Feng et al., 2020; Ye et al., 2021; Wang et al., 2021b; Guo et al., 2021; Wang et al., 2022; Li et al., 2022c) and using vector distances to quantify similarity. We use GraphCodeBERT (Guo et al., 2021) as the base similarity model since its model checkpoint is off-the-shelf, widely used, and has been pre-trained on large-scale code.
To enhance the function-level representation of GraphCodeBERT further from a pre-trained transformer, we employ self-supervised constrastive learning. Following (Gao et al., 2021), we adapt the unsupervised SimCSE method to the code domain, using standard dropout as the data augmentation method for contrastive learning. For an input code snippet $x$, we use the last-layer hidden states of [CLS] token as code representation. During the unsupervised training stage, we add an MLP layer to obtain the final representation $\mathbf{h}$ of $x$.
@@
Where $\tau$ is a temperature hyperparameter and is set to $0.1$. The $sim(\mathbf{h}_{i},\mathbf{h}^{\prime}_{i})$ is the cosine similarity $\frac{\mathbf{h}_{i}\cdot\mathbf{h}^{\prime}_{i}}{|\mathbf{h}_{i}||\mathbf{h}^{\prime}_{i}|}$.
The MLP layer is dropped after the unsupervised contrastive learning stage, and we only use the last-layer hidden states of the [CLS] token as code snippet representation following (Gao et al., 2021). For a given code snippet $x$ and its rewriting $x^{\prime}$ sampled from $\mathcal{G}$ as introduced in Section 3.3, we obtain the final representations of $x$ and $x^{\prime}$ by feeding them into the similarity model $\mathcal{M}$. It is worth noting that our framework is not dependent on a specific similarity model; $\mathcal{M}$ can be any implementation that accurately model code similarity, such as OpenAI’s text embedding services (Greene et al., 2022).
@@
A larger $m$ can lead to a more accurate estimation of expectation but at the cost of generating more rewritten code. Excitingly, our experiments found that only using 4 rewrites can achieve outstanding detection performance.
@@ 4.1. Benchmarks
Considering the absence of existing benchmarks for evaluating synthetic code detectors, we build two synthetic code detection benchmarks in Python using APPS (Hendrycks et al., 2021) and MBPP (Austin et al., 2021). The choice of Python as the programming language for constructing our benchmark stems from several key considerations. Firstly, a significant portion of training data for prevalent code generation models like CodeGen (Nijkamp et al., 2022), Incoder (Fried et al., 2022), and StarCoder (Li et al., 2023) is constituted by Python code. These models exhibit notably superior code generation capabilities in Python than in other languages. Consequently, the challenge and significance of detecting code generated by LLMs are most pronounced within the context of Python. Secondly, popular datasets used for code generation evaluation, such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and APPS (Hendrycks et al., 2021), primarily provide Python test cases. We adopted Python as our chosen language to align with this prevailing practice. Furthermore, Python ranks among the most widely used programming languages, adding practical significance to validating our detection performance. Specifically, for the generation of synthetic code, we utilize the natural language description of each sample as a prompt and employ four widely-used code generation tools (CodeLlama (Rozière et al., 2023), StarChat (Tunstall et al., 2023), GPT-3.5 (OpenAi, 2022) and GPT-4 (OpenAI, 2023)).
@@
We select two open-sourced Code LLMs, CodeLlama-13B-Instruct (Rozière et al., 2023) and StarChat-Alpha (Tunstall et al., 2023), along with two proprietary generation APIs, GPT-3.5-turbo and GPT-4 (OpenAI, 2023). According to the survey (StackOverflow, 2023) conducted by StackOverflow, GPT-3.5-turbo is the most extensively used tool in daily programming. GPT-4 represents the state-of-the-art in code generation tools to date. Consequently, we have chosen these two models to typify proprietary generation APIs. CodeLlama (Rozière et al., 2023) and StarChat (Tunstall et al., 2023) are two potent and popular open-sourced code generation tools with millions of downloads on HuggingFace [^3] https://huggingface.co/codellama [^4] https://huggingface.co/HuggingFaceH4/starchat-alpha. We opted for the popular 13B and 15.5B versions due to resource constraints.
@@
APPS (Hendrycks et al., 2021) is a benchmark for code generation. Each sample in APPS is a coding problem taken from coding contest websites, e.g., CodeForces, LeetCode, and HackerRank, and contains problem descriptions and submitted solutions for the problem. The original APPS Benchmark comprises 5,000 test data and 5,000 training data instances. In our construction, we exclusively utilized the test data. The rationale for excluding the training data stems from concerns that the training data might have been incorporated into the training data of code LLMs. If the training data are used as prompts and those prompts are present in the LLM’s training data, it could potentially result in code generation that perfectly matches human-written code, leading to inaccuracies in evaluating the detection outcomes. Within the pool of 5,000 test data instances, we identified and removed prompts with web hyperlinks from the Mdarkdown format of coding challenges, as such links could affect code quality and are rarely used by developers in practice. This reduced the dataset to 3,846 instances. We then randomly selected 1,540 instances (40%) from this refined set and used 770 of them for synthetic code generation.
For those 770 problems, we generate synthetic code using the chain of thought prompting (Wei et al., 2022b) to improve the generation quality using four generation tools. The exact prompt used is shown in Figure 4. The generation temperature is set to $0.7$, and $top\_p$ is set to 0.95. We removed all the comments and empty lines for the generated code. For the remaining 770 problems (1540 minus 770), we randomly sample a solution from all valid solutions as human-written code. The detailed dataset statistics are listed in Table 1.
@@
Table 1. APPS Benchmark Statistics. We count the averaged number of chars and lines for human-written/synthetic code for all versions of benchmark generated by four different code generation tools.
@@
MBPP (Austin et al., 2021) is another benchmark for evaluating AI code generation. It contains 1,000 crowd-sourced Python programming problems designed to be solvable by entry-level programmers, covering programming fundamentals, standard library functionality, etc. Compared to APPS, MBPP’s problems are more straightforward and frequently encountered in daily programming practice. We follow the same construction process as APPS. The detailed dataset statistics are listed in Table 2.
@@ 4.2. Similarity Model Training Details
We continue to train a code similarity model by exploiting unsupervised SimCSE from initial GraphCodeBERT as introduced in Section 3.4. In the SimCSE training stage, we collect 160k code snippets from CodeSearchNet (Husain et al., 2020) Python subset as training data. The GraphCodeBERT model trained by SimCSE is fixed for all subsequent experiments once training is complete. We use the Adam optimizer and set the maximum learning rate to $41e-$ 4 with linear decay. We use 4 NVIDIA 3090s with batch size 16 on each GPU and train the model for 5 epochs. In addition to the GraphCodeBERT model, we also experimented with the UniXcoder (Guo et al., 2022) as another similarity model by following the same SimCSE training process.
@@ 4.3. Detector LLMs
Our approach necessitates a rewriting LLM, $\mathcal{G}$, to function as the detector LLM. Additionally, other zero-shot detectors for text, such as GLTR and DetectGPT, require a scoring LLM for likelihood estimation, as introduced in Section 2.2. We examine CodeLlama, StarChat-Alpha, and GPT-3.5 as detector LLMs for our method while using CodeLlama and StarChat-Alpha for baseline comparison. We employ nucleus sampling for code rewriting, setting the $top\_p$ to $0.95$ and the temperature to $0.8$.
@@ 4.5. Baselines
We consider the following zero-shot detection methods utilizing a surrogate model to approximate the true distribution of the generative model: $\mathbf{\log p(x)}$ (Gehrmann et al., 2019), LogRank, Rank, Entropy (Mitchell et al., 2023), LRR, NPR (Su et al., 2023) and DetectGPT (Mitchell et al., 2023). To adapt these detection methods to code content and effectively approximate the true distribution of the generative code LLM, the surrogate score model should be replaced with LLMs trained on large-scale code content. In this study, we examine two open-sourced code LLMs: StarChat-Alpha and CodeLlama. We also select two supervised detectors and compare them to the zero-shot detectors: GPTZero (GPTZero, 2023), a leading AI content detection services, is trained on millions of synthetic texts sampled from various generative models, including ChatGPT, GPT-4 and Bard. OpenAI-Detector (OpenAI, 2019), an open-sourced detector, is trained on texts sampled from GPT-2 and is based on Roberta-large (Liu et al., 2019).
@@
DetectGPT (Mitchell et al., 2023) uses the average token probability disparity between the given text and perturbed texts to detect whether the given text is located at the probability curvature of the generative model. To better adapt DetectGPT for the code domain, we made minor modifications. Specifically, we replace the original perturbation model, T5-large (Raffel et al., 2020) with CodeT5-large (Wang et al., 2021a). All other settings are the same as in the original implementation.
@@ 5. Experiment Results
Table 3. Main Results. The first two rows list two benchmarks and their corresponding four Generation Tools. The following sections detail the AUROC scores of our methods, with variations in $m$, alongside seven other zero-shot detectors, using three Detector LLMs. CodeLlama abbreviates CodeLlama-13B-Instruct, StarChat representes StarChat-Alpha-15B, and GPT-3.5 refers to GPT-3.5-turbo. Red indicates the best performance, while blue indicates the best performance among all seven baselines.
@@ 5.1. Main Results
We list the AUROC score of all zero-shot detectors on the APPS and MBPP benchmarks in Table 3. Among the seven baselines, LogRank, LRR, and NPR are the most compelling methods. However, the detection performance of baseline zero-shot detectors, including $\log p(x)$, DetecGPT, LRR, and NPR drop significantly on the code benchmarks compared to their reported performance on detecting general texts (Mitchell et al., 2023; Su et al., 2023). This observation holds true regardless of the generation tools and detector LLMs used. The main reason, as mentioned earlier, lies in the fact that these SOTA zero-shot text detectors are based on the log probability of tokens statistically. In the code domain, code exhibits a more uniform grammatical structure, often with a relatively fixed space for the next code tokens. As a result, there are many ”low-entropy” tokens in code, which significantly reduces the number of effective tokens that SOTA text detection methods can use to differentiate between human-written code and synthetic code, as illustrated in Figure 1. Furthermore, We observe that the previous baselines demonstrate improved performance when the Detector LLM and the Generation LLM are identical, and exhibit a decline in performance when they do not match.
As shown in Table 3, our proposed methods significantly outperform previous methods by all four prevalent code generation tools, with a 20.5% improvement on the APPS benchmark and 29.1% improvement on the MBPP benchmark. This improvement is primarily due to our proposed methods of detecting synthetic code from a more holistic perspective instead of relying on token-wise scores. The holistic code rewriting and similarity measurement approach effectively avoids the inherent shortcomings of previous methods when applied to the code domain, while also leveraging the LLM’s consistent coding style and the tendency of LLMs to output the most confident and likelihood code during generation.